### v2

# ANPrompt: Anti-noise Prompt Tuning for Vision-Language Models

Abstract
Prompt tuning has emerged as an efficient and effective technique for adapting vision-language models (VLMs) with low computational overhead. However, existing methods often overlook the vulnerability of prompt-tuned VLMs to weak semantic perturbations—such as subtle image or text noise—that degrade their generalization to unseen classes. To address this limitation, we propose ANPrompt, a novel prompt tuning framework designed to enhance robustness under such perturbations. ANPrompt first constructs weak noise text features by fusing original and noise-perturbed text embeddings, which are then clustered to form noise prompts. These noise prompts are integrated with learnable prompt tokens to generate anti-noise prompts, which are injected into the deeper layers of both image and text encoders. To further capture the noise-aware visual semantics, ANPrompt computes the Noise-Resistant Visual Prompt Prototype (NRVPP) by averaging the output prompt tokens from the vision encoder. Finally, ANPrompt introduces alignment, robustness, and anti-noise objectives by computing a Weak semantic noise Alignment Loss (WALoss) alongside the standard cross-entropy and sim loss. Experiments across 11 benchmarks demonstrate that ANPrompt consistently outperforms existing prompt tuning approaches, achieving superior robustness to semantic noise and improved generalization to novel categories.
Introduction
Vision-language models (VLMs) like CLIP have demonstrated impressive performance on open-vocabulary recognition tasks  (Hua et al. 2025; Weng et al. 2023; Xu et al. 2024), but adapting such large models to downstream tasks via full fine-tuning can be computationally prohibitive. To address this, prompt tuning (Zhu et al. 2024; Lafon et al. 2024; Shi et al. 2024) has emerged as a lightweight alternative by optimizing only a small set of learnable tokens, achieving state-of-the-art results across various domains. Despite its efficiency, early prompt tuning methods such as CoOp (Zhou et al. 2022b) often overfit to seen classes, resulting in limited generalization to novel categories and vulnerability to subtle semantic perturbations.
To enhance transferability, recent approaches have integrated inductive biases into the prompting process. For instance, CoCoOp conditions prompts on image features to mitigate overfitting; KgCoOp constrains semantic drift by leveraging frozen text embeddings; and MaPLe adopts multimodal prompts to achieve more robust alignment. More advanced techniques such as PromptSRC and CoPrompt regularize prompt learning using frozen zero-shot logits, while MMRL and TAC construct shared or clustered semantic spaces. Although designed to reduce overfitting, consistency constraints based on frozen features may introduce noise information such as misleading text or visual clutter caused by domain gaps, ultimately weakening the robustness of Vision-Language Models.
Figure 1: Weak semantic perturbations in image and text: CAMs may highlight irrelevant regions (e.g., background or co-occurring objects), while subtle textual cues (e.g., “dark stripes”) may mislead class semantics (e.g., shift from “cat” to “tiger”).
Figure 2: Different with existing methods use frozen logits as constraint, risking overfitting under weak semantic perturbations. ANPrompt introduces weak semantic perturbations to generate soft logits and anti-prompt to improves generalization.
As shown in Figure 1, we define subtle noise in images and texts as weak semantic perturbations, which appear as either misleading text (e.g., “a cat resembling a small tiger”) or visual noise (e.g., background clutter or co-occurring objects). While high-quality image-text pairs exhibit clear object focus and accurate descriptions, existing prompt tuning methods often struggle to handle such subtle shifts, leading to significant performance drops when adapting to novel categories. This highlights the urgent need for VLMs capable of learning robust, noise-resilient cross-modal representations to maintain alignment consistency and improve generalization.
To address the issue of weak semantic perturbations, we propose ANPrompt — a prompt tuning framework specifically designed to enhance robustness under such perturbations. As illustrated in Figure 2, existing methods rely on frozen feature constraints and thus tend to suffer from overfitting when exposed to weak semantic perturbations. In contrast, our ANPrompt framework intentionally incorporates weak semantic perturbation information into prompt tokens and logits, thereby boosting the robustness of both the learned prompts and class distributions in scenarios involving weak semantic perturbations. Specifically, ANPrompt randomly samples two intra-class descriptions from a language cache: a main text that conveys core semantics, and a noise text that introduces minor linguistic variations. Their fusion generates a Weak Noise Frozen Text Feature, which incorporates weak semantic perturbation information while preserving class identity. These features are clustered into noise prompts and injected as biases into learnable tokens, forming Anti-noise Prompts. These prompts are then inserted into the higher layers of both text and vision encoders to introduce weak semantic perturbation information at the prompt representation level. After encoding, ANPrompt extracts image Feature, text Feature, and the Noise-Resistant Visual Prompt Prototype (NRVPP). The latter is obtained by mean-pooling the visual prompt tokens from the final layer of the visual encoder. Since Anti-noise Prompts are injected into deeper encoder layers, the resulting NRVPP inherit weak semantic perturbations.
To stabilize training under both weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss (WALoss) between Weak Noise Logits and Logits set, which enforces alignment consistency under controlled weak semantic perturbations. Additionally, we employ a CE and sim loss to enable the model to learn task-relevant knowledge under weak semantic perturbations. By combining these losses, ANPrompt can improve the robustness of VLMs under weak semantic perturbations.
Extensive experiments conducted on 11 benchmarks and base-to-new splits show that ANPrompt significantly outperforms existing prompt tuning methods, achieving state-of-the-art performance in terms of robustness and generalization.
Our main contributions are summarized as follows:
  • We propose ANPrompt, a prompt tuning framework that improve the robustness of VLMs by injecting weak semantic perturbations into both prompt tokens and logits pathways.
  • We introduce a Weak semantic noise Alignment Loss (WALoss) to enforce alignment consistency under controlled weak semantic perturbations.
  • We develop a weak semantic perturbations prompt injection mechanism that clusters weak noise frozen text features into noise prompts and injects them into learnable tokens to get anti-noise prompt, which is used to inject weak semantic perturbations at the prompt representation level.
  • ANPrompt achieves state-of-the-art performance on 11 benchmark datasets under base-to-new generalization settings.
Related Work
Vision-Language Models
Pretrained vision-language models (VLMs) such as CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021) achieve strong open-vocabulary and cross-modal capabilities through contrastive learning on large-scale image-text pairs. Building on this foundation, models like BLIP (Li et al. 2022), FILIP (Yao et al. 2021), Florence (Yuan et al. 2021), and LiT (Zhai et al. 2022) enhance fine-grained alignment and data efficiency via retrieval augmentation(Li et al. 2024a), self-supervised learning, and stronger backbones. NLIP (Huang et al. 2023) further improves robustness by introducing noise-harmonization based on alignment uncertainty and noise-completion via concept-guided caption generation. These advances enable effective zero-shot and few-shot transfer for tasks such as retrieval, classification, and grounding. However, adapting such large models to downstream tasks (Li et al. 2025a, 2023) remains costly due to domain shifts and the high overhead of full fine-tuning. To address this, parameter-efficient tuning methods—such as adapter-based approaches (Gao et al. 2024; Yang et al. 2024b; Song et al. 2023) and prompt learning (Zhou et al. 2022b, a; Khattak et al. 2024)—have been proposed to enable lightweight adaptation without retraining the entire model.
Prompt Learning
Prompt learning has become an efficient strategy to adapt VLMs without full fine-tuning. Early methods such as CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a) explore soft and image-conditioned prompts to enhance generalization. Subsequent works improve robustness and transferability through multimodal regularization (MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b)), gradient alignment (ProGrad (Zhu et al. 2023)), hierarchical and knowledge-guided prompts (HPT (Wang et al. 2024), TCP (Yao, Zhang, and Xu 2024)), and distillation-based initialization (CLIP-kd (Yang et al. 2024a), ComKD-CLIP (Chen et al. 2024), PromptKD (Li et al. 2024b), CasPL (Wu et al. 2024)). Other methods focus on attribution-based prompting (ATPrompt (Li et al. 2024c)), two-stage decoupled optimization (DPC (Li et al. 2025b), DePT (Zhang et al. 2024), 2SFS (Farina et al. 2025)), or cross-modal adapters (MMA (Yang et al. 2024b)) for few-shot settings. Consistency learning (CoPrompt (Roy and Etemad 2024)) and task-aware clustering (TAC (Hao et al. 2025)) further enhance adaptability. However, these approaches largely target generalization and efficiency, with limited attention to robustness against weak semantic perturbations.
Prompting under Weak Semantic Perturbation Noise
Recent works have explored robust prompt learning under noisy conditions. Methods such as NLPrompt (Pan et al. 2025), JoAPR (Guo and Gu 2024), and ArGue (Tian et al. 2024) mitigate label noise through loss design, sample filtering, and LLM-generated attributes, while ClipCleaner (Feng, Tzimiropoulos, and Patras 2024) leverages CLIP-based sample selection to decouple training from noise filtering. However, these approaches mainly focus on filtering noise, which risks discarding informative variations and often limits generalization. In contrast, we propose ANPrompt, which embraces weak semantic perturbations as supervision signals. By integrating perturbed prompts with clustering-based bias representations, ANPrompt enables robust and discriminative prompt learning without relying on brittle filtering heuristics.
Figure 3: ANPrompt randomly samples a main and a noise sentence from the same class to construct the Weak Noise Frozen Text Feature, which is used to generate Anti-noise Prompts. These prompts are injected into the encoder’s deep layers. During encoding, we extract the Prompted Image/Text Features and the Noise-Resistant Visual Prompt Prototype (NRVPP) to compute four types of logits, which are supervised by WALoss and auxiliary losses to improve robustness against weak semantic perturbations.
Methodology
Preliminary
Vision-language models (VLMs) such as CLIP (Radford et al. 2021) learn a joint embedding space for images and texts through contrastive learning over large-scale image-text pairs. These models exhibit strong zero-shot generalization but are highly sensitive to semantic variations in either modality.
A typical VLM consists of a vision encoder and a text encoder. The vision encoder tokenizes the input image into patches, encodes them through Transformer layers, and outputs a class token representing global semantics. The text encoder processes class-dependent prompts (e.g., “a photo of a [CLASS]”), encodes them with Transformers, and uses the [EOS] token as the prompt embedding.
The classification score for class $c$ is computed via cosine similarity:
$p(y=c\mid f)=\frac{\exp(\text{sim}(f,w_{c})/\tau)}{\sum_{i=1}^{C}\exp(\text{sim}(f,w_{i})/\tau)}$ (1)
where $f$ is the image feature, $w_{c}$ is the text feature for class $c$, $\tau$ is a learnable temperature, and $\text{sim}(\cdot)$ denotes cosine similarity.
Overview
We propose ANPrompt — a robust prompt tuning framework that injects weak semantic perturbations into the prompting and alignment processes to enhance the robustness of VLMs under such perturbations. Specifically, ANPrompt simulates weak semantic perturbations by fusing two intra-class textual descriptions from the language cache, generating a Weak Noise Frozen Text Feature. These Weak Noise Frozen Text Features are clustered to obtain noise prompts, which are then projected and added to learnable prompt tokens to form Anti-noise Prompts. These prompts are inserted into the deeper layers of both vision and text encoders, enabling ANPrompt to inject weak semantic perturbations at the prompt level. Subsequently, prompted image features, prompted text features, and NRVPP can be obtained for computing alignment logits, robustness logits, weak noise logits, and anti-noise logits. During training, ANPrompt combines the Weak Semantic Noise Alignment Loss (WALoss), cross-entropy loss, and SimLoss to enhance the robustness of VLMs under weak semantic perturbations.
ANPrompt Framework
As shown in Figure 3. The ANPrompt framework proceeds through four sequential stages: (1) Weak Noise Frozen Text Feature Construction, (2)Anti-noise Prompt Generation, (3) Feature Encoding, and (4) Logit Computation.
1. Weak Noise Frozen Text Feature Generation: ANPrompt samples two intra-class descriptions: a main text that encodes core semantics, and a noise text that introduces subtle phrasing variations (e.g., attribute reordering). These two texts are encoded by the frozen CLIP text encoder into $f_{m}$ and $f_{n}$ respectively, and fused as follows:
$f_{w}=f_{m}+\alpha f_{n}$ (2)
where $\alpha$ regulates the perturbation strength. Since the Weak Noise Frozen Text Feature $f_{w}$ incorporates contextual information from different sentences, it inevitably captures contextual cues that lean toward semantically adjacent yet incorrect classes. Consequently, weak noise text features inherently encode weak semantic perturbations information.
2. Anti-noise Prompt Generation: To map weak semantic perturbation information into the prompt space, ANPrompt performs K-means clustering on all $f_{w}$ to extract representative semantic centers, which are referred to as noise prompts. These noise prompts are added to the learnable prompt tokens $P_{c}$ to form intermediate anti-noise prompts $P_{a}$. These intermediate prompts are then fed into two separate fully connected layers to generate modality-specific Anti-noise Prompts, which is injected into the deeper layers of both encoders.
3. Feature Encoding. After prompt injection, we extract three types of features: the Prompted Image Feature $f_{v}$, encoded from the final image class token; the Prompted Text Feature $f_{t}$, derived from the final [EOS] token; and the Noise-Resistant Visual Prompt Prototype (NRVPP) $f_{N}$, obtained by mean-pooling the visual Anti-noise Prompts after encoding, which captures weak semantic perturbations information.
4. Logit Computation. We compute four sets of logits: the Alignment Logits $\ell_{A}=f_{v}\cdot f_{t}^{\top}$, which measure the similarity between the prompted image and text features; the Robustness Logits $\ell_{R}=f_{v}\cdot f_{w}^{\top}$, which assess the model’s robustness against weak semantic noise in text; the Weak Noise Logits $\ell_{W}=f_{N}\cdot f_{w}^{\top}$, which reflect the alignment between noise-resistant visual features and weakly perturbed text features; and the Anti-noise Logits $\ell_{\text{final}}=\theta\ell_{A}+(1-\theta)\ell_{R}$, which integrate alignment and robustness for final prediction.
To stabilize training under weak semantic perturbations, we apply $\mathcal{L}_{\text{WA}}$ between $\ell_{W}$ and $\ell_{A}$ for alignment consistency under weak semantic perturbations, $\mathcal{L}_{\text{CE}}$ on $\ell_{\text{final}}$ for task supervision, and $\mathcal{L}_{\text{sim}}$ between frozen and prompted features to enhance semantic consistency.
The final loss is:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{sim}}+\gamma\mathcal{L}_{\text{WA}}$ (3)
where $\mathcal{L}_{\text{CE}}$ denotes a cross-entropy loss, $\mathcal{L}_{\text{sim}}$ represents a cosine loss, and $\mathcal{L}_{\text{WA}}$ is defined as follows. $\lambda$ serves as the weight of $\mathcal{L}_{\text{sim}}$, while $\gamma$ is a variance-adaptive weight used to regulate the strength of weak semantic perturbations.
Weak Noise Alignment Loss
To enforce alignment consistency under weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss ($\mathcal{L}_{\text{WA}}$) that minimizes the divergence between $\ell_{W}$ and $\ell_{A}$. Since KL divergence measures the discrepancy between distributions, minimizing it encourages the perturbed prediction $\ell_{W}$ to stay close to the original alignment $\ell_{A}$, penalizing semantic shifts and improving robustness while preserving alignment structure.
To compute $\mathcal{L}_{\text{WA}}$, we first derive $\ell_{A}$ from the prompted image and text features. Then, to obtain $\ell_{W}$, we compute the noise-resistant visual prototype $f_{N}$ by averaging the visual prompt token features $f_{v}^{o}\in\mathbb{R}^{N\times C}$:
$f_{N}=\frac{1}{N}\sum_{i=1}^{N}f_{v}^{o(i)}$ (4)
This averaging suppresses token-level variations and mitigates visual semantic distraction, especially under anti-prompts with weak semantic noise, where individual tokens may attend to irrelevant regions such as backgrounds or co-occurring objects. The weak noise logits are then computed as:
$\ell_{W}=f_{N}\cdot f_{w}^{\top}$ (5)
Finally, the weak alignment loss is defined as:
$\mathcal{L}_{\text{WA}}=\gamma\cdot\text{KL}(\text{softmax}(\ell_{R})\|\text{softmax}(\ell_{W}))$ (6)
with the variance-adaptive weight $\gamma$ given by:
$\gamma=\frac{1}{\text{std}(\ell_{R})\cdot|\ell_{R}|+\epsilon_{0}}$ (7)
where $\text{std}(\ell_{R})$ denotes the standard deviation over all elements in $\ell_{R}$, $|\ell_{R}|$ represents the total number of elements (i.e., numel() of logits), and $\epsilon_{0}$ is a small constant to prevent division by zero. This global scaling dynamically adjusts the loss weight based on the dispersion and size of the logits, stabilizing optimization under varied prediction confidence levels.
Weak Semantic Perturbations Prompt Injection Mechanism
We introduce a prompt injection mechanism based on weak semantic perturbations. Specifically, we first cluster the weak noise frozen text features $f_{w}\in\mathbb{R}^{N\times C}$ into $K$ noise prompts $P_{w}\in\mathbb{R}^{k\times C}$ using K-Means, where $k$ matches the number of learnable prompt tokens $P_{c}\in\mathbb{R}^{k\times C}$ to ensure one-to-one integration.
These noise prompts are added to the learnable tokens to form the anti-noise prompt:
$P_{a}=P_{c}+\epsilon P_{w}$ (8)
where $\epsilon$ controls the strength of injected weak perturbations. The anti-noise prompt $P_{a}$ is then projected into modality-specific spaces via a mapping layer.
Finally, the projected prompts $P_{a}^{\text{image}}$ and $P_{a}^{\text{text}}$ are injected into the visual and textual input sequences:
$\{P_{a}^{\text{image}},e_{\text{cls}},e_{1},e_{2},\dots,e_{M}\}$ (9)
$\{t_{\text{SOS}},P_{a}^{\text{text}},t_{1},t_{2},\dots,t_{L},c_{k},t_{\text{EOS}}\}$ (10)
This injection enriches the prompt space with weakly perturbed semantic cues, promoting robustness against subtle noise while maintaining structural alignment with the original prompt format.
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bold. We can see that our method achieves the best performance compared with other state-of-the-art methods.
Method
Ref.
Average ImageNet Caltech101 OxfordPets
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CLIP ICML22 69.34 74.22 71.70 72.43 68.14 70.22 96.84 94.00 95.40 91.17 97.26 94.12
CoOp IJCV22 82.69 63.22 71.66 76.47 67.88 71.92 98.00 89.81 94.20 93.67 95.29 94.47
CoCoOp CVPR22 80.47 71.69 75.83 77.60 70.75 74.00 97.96 93.81 95.84 95.20 97.69 96.43
MaPLe CVPR23 82.28 75.14 78.55 76.66 70.54 73.47 97.74 94.36 96.02 95.43 97.76 96.58
PromptSRC ICCV23 84.26 76.10 79.97 77.60 70.73 74.01 98.10 94.03 96.02 95.33 97.30 96.30
CoPrompt ICLR24 84.00 77.23 80.48 77.67 71.27 74.33 98.27 94.90 96.55 95.87 98.10 96.87
TextRefiner AAAI25 79.74 74.32 76.94 76.84 70.54 73.56 98.13 94.43 96.24 95.27 97.65 96.45
MMRL CVPR25 85.68 77.16 81.20 77.90 71.30 74.45 98.97 94.50 96.68 95.90 97.60 96.74
ANprompt 86.15 77.70 81.70 77.83 71.17 74.35 98.97 94.73 96.80 95.73 97.17 96.44
Method
Ref.
StandfordCars Flowers102 Food101 FGVCAircraft
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CLIP ICML22 63.37 74.89 68.65 72.08 77.80 74.83 90.10 91.22 90.66 27.19 36.29 31.09
CoOp IJCV22 78.12 60.40 68.13 97.60 59.67 74.06 88.33 82.26 85.19 40.44 22.30 28.75
CoCoOp CVPR22 70.49 73.59 72.01 94.87 71.75 81.71 90.70 91.29 90.99 33.41 23.71 27.74
MaPLe CVPR23 72.94 74.00 73.47 95.92 72.46 82.56 90.71 92.05 91.38 37.44 35.61 36.50
PromptSRC ICCV23 78.27 74.97 76.58 98.07 76.50 85.95 90.67 91.53 91.10 42.73 37.87 40.15
CoPrompt ICLR24 76.97 74.40 75.66 97.27 76.60 85.71 90.73 92.07 91.40 40.20 39.33 39.76
TextRefiner AAAI25 71.40 70.90 71.15 95.92 74.33 83.76 90.88 91.43 91.15 35.35 35.87 35.61
MMRL CVPR25 81.30 75.07 78.06 98.97 77.27 86.78 90.57 91.50 91.03 46.30 37.03 41.15
ANprompt 83.57 74.63 78.85 98.60 77.30 86.66 90.63 91.50 91.06 49.67 36.60 42.14
Method
Ref.
SUN397 DTD EuroSAT UCF101
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CLIP ICML22 69.36 75.35 72.23 53.24 59.90 56.37 56.48 64.05 60.03 70.53 77.50 73.85
CoOp IJCV22 80.60 65.89 72.51 79.44 41.18 54.24 92.19 54.74 68.69 84.69 56.05 67.46
CoCoOp CVPR22 79.74 76.86 78.27 77.01 56.00 64.85 87.49 60.04 71.21 82.33 73.45 77.64
MaPLe CVPR23 80.82 78.70 79.75 80.36 59.18 68.16 94.07 73.23 82.35 83.00 78.66 80.77
PromptSRC ICCV23 82.67 78.47 80.52 83.37 62.97 71.75 92.90 73.90 82.30 87.10 78.80 82.74
CoPrompt ICLR24 82.63 80.03 81.31 83.13 64.73 72.79 94.60 78.57 85.84 86.90 79.57 83.07
TextRefiner AAAI25 80.96 76.49 78.66 75.35 58.09 65.60 74.57 75.82 73.68 82.52 75.01 78.59
MMRL CVPR25 83.20 79.30 81.20 85.67 65.00 73.82 95.60 80.17 87.21 88.10 80.07 83.39
ANprompt 83.07 79.07 81.02 85.20 65.10 73.80 95.53 87.33 91.21 88.80 80.07 84.21
Table 2: Results from cross-dataset evaluation. ANprompt provides the highest average accuracy.
Source Target
ImageNet Caltech Pets Cars Flowers Food Aircraft SUN397 DTD EuroSat UCF101 Average
CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.90 46.39 66.55 63.88
CoCoOp 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
MaPLe 70.72 93.53 90.55 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.06 66.10
PromptSRC 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
CoPrompt 70.80 94.50 90.75 65.67 72.70 86.43 20.40 67.57 47.07 51.90 69.73 67.00
TAC 72.77 94.53 90.67 65.30 72.20 85.83 23.53 67.63 47.57 48.07 70.00 66.53
ANprompt 71.13 94.70 91.00 66.00 72.87 86.23 26.10 67.47 45.83 51.97 69.27 67.14
Experiments
Evaluation Task
Base-to-Novel Generalization
To assess the zero-shot generalization capability of ANprompt, we divide each dataset into base and novel categories. The model is trained on the base categories using a few-shot setting, and its performance is evaluated on both the base and novel categories without additional training.
Cross-Dataset Evaluation
To examine the transferability of the model across datasets, we conduct experiments where the model is trained on a few-shot subset of the ImageNet dataset and directly evaluated on other datasets, without any further adaptation or fine-tuning.
Domain Generalization
To test the model’s resilience to domain shift, we use the ImageNet-trained model and evaluate it on four distinct ImageNet-derived datasets that introduce various domain differences, allowing us to measure performance under out-of-distribution scenarios.
Evaluation Settings
Datasets
In the tasks of Base-to-New Generalization and Cross-Dataset Evaluation, we followed the experimental setup of CoOp (Zhou et al. 2022b), and evaluated on 11 public datasets, including ImageNet, Caltech (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al. 2012), Flowers (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Gool 2014), StanfordCars (Krause et al. 2013), FGVCAircraft (Maji et al. 2013), EuroSAT (Helber et al. 2019), UCF101 (Kay et al. 2017), DTD (Cimpoi et al. 2014) and SUN397 (Xiao et al. 2010). We also use ImageNet (Deng et al. 2009) as the source dataset, and use its four variants ImageNetV2 (Recht et al. 2019), ImageNet-Sketch (Wang et al. 2019), ImageNet-A (Hendrycks et al. 2021b) and ImageNet-R (Hendrycks et al. 2021a) as the target dataset to evaluate the domain generalization task.
Implementation Details
We set the learning rate to 0.001 and adopt the Adam optimizer to train ANPrompt. The model is trained for 10 epochs with a batch size of 4, using ViT-B/16 as the vision backbone. For the LLM cache, we incorporate prompts from CoPrompt (Roy and Etemad 2024) and HPT (Wang et al. 2024). And we set the $\theta$ and $\alpha$ with 0.7 and 0.001. All experiments are conducted on a single NVIDIA RTX 3090 Ti GPU. Reported results are averaged over three random seeds.
Experimental Performance
We compare the results of our method on three evaluation tasks with other advanced prompt learning methods, including zero-shot CLIP and some prompt learning methods, such as CoOp (Zhou et al. 2022b), CoCoOp (Zhou et al. 2022a), MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), CoPrompt (Roy and Etemad 2024), TextRefiner (Xie et al. 2025) and MMRL (Guo and Gu 2025). We also added some targeted ablation experiments to prove the effectiveness of the proposed ANPrompt method.
Base-to-New Generalization
Table 1 shows the performance of ANPrompt and other comparison methods on B2N generalization tasks. Compared with the most advanced MMRL method, ANPrompt improves the base class by 0.44% and the new class by 0.43%. Combining the base class and the new class, ANPrompt achieves a performance improvement of 0.44% on the harmonic average. It is worth noting that compared with MMRL (Guo and Gu 2025), ANPrompt achieves significant performance improvement on EuroSAT dataset: 6.03% on new class and 3.31% on harmonic average. By transforming weak semantic noise into constructive training signals, ANPrompt achieves performance improvement by fusing complementary semantic perturbation sentences.
Cross-Dataset Evaluation
Table 2 shows the performance comparison between our ANPrompt and existing methods in cross-dataset evaluation. On the ImageNet source dataset, ANPrompt shows comparable performance to other competitive methods. On the target domain dataset, ANPrompt achieved the best performance on six of the ten datasets, showing excellent generalization ability, especially on fine-grained datasets, such as FGVCAircraft and EuroSAT. Specifically, with an average of 10 datasets, ANPrompt achieved the highest accuracy rate of 67.14%. Meanwhile, compared with the suboptimal method CoPrompt, ANPrompt achieved 0.14% performance improvement. Different from the existing methods, we use the anti-noise visual prompt prototype to align the weak noise frozen text features to generate soft, context-aware supervision signals, thus improving cross-domain performance.
Table 3: Results on domain generalization task.
Source Target
ImageNet -V2 -Sk -A -R Average
CLIP 66.73 60.83 46.15 47.77 73.96 57.18
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.90
MaPLe 70.72 64.07 49.15 50.90 76.98 60.26
PromptSRC 71.27 64.35 49.55 50.90 77.80 60.63
CoPrompt 70.80 64.25 49.43 50.50 77.51 60.42
MMRL 72.03 64.47 49.17 51.20 77.53 60.59
ANprompt 71.13 64.63 49.13 50.37 77.47 60.40
Domain Generalization
We compare the performance of our ANPrompt with that of six previous methods on cross-domain data sets. As shown in Table 3, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability. Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
Ablation Experiments
Componment Ablation
We conduct an ablation study to evaluate the contribution of each component in ANPrompt, including the Weak noise Frozen Text feature (TextNoise), Weak Alignment Loss (WALoss), and Anti-noise Prompt injection. As shown in Table 4, introducing WALoss or Anti-Prompt individually yields consistent improvements over the baseline, with gains of +0.06 and +0.19 respectively. Notably, the addition of TextNoise leads to further gains when combined with either WALoss (+0.34) or Anti-Prompt (+0.40). When all three components are enabled, ANPrompt achieves the highest harmonic mean (HM) of 81.70, surpassing the baseline by +0.79. These results verify the complementary nature of each module and their collective effectiveness in enhancing robustness under noise.
Table 4: The Effectiveness of different components in ANPrompt. The TextNoise stands for the Weak noise Frozen Text feature.
TextNoise WALoss Anti-Prompt HM
80.91
80.97(+0.06)
81.10(+0.19)
81.25(+0.34)
81.27(+0.36)
81.31(+0.40)
81.70(+0.79)
Effect of Prompt Token Length.
We study the effect of varying the number of prompt tokens $T$ on performance. As shown in Table 5, increasing $T$ from 1 to 5 gradually improves performance across base and novel classes, leading to a peak harmonic mean (HM) of 81.70% at $T=5$. This suggests that moderate prompt length enhances generalization. However, performance saturates at $T=6$, where no further gain is observed, which proved the effectiveness of the ablation experiment on token length. Therefore, setting $T=5$ offers the best trade-off between expressiveness and stability.
Table 5: The length of Prompt Token.
$T$ 1 2 3 4 5 6
$B$ 85.80 86.08 86.13 86.08 86.15 86.09
NN 75.87 75.25 75.47 75.87 77.70 77.67
HMHM 80.53 80.30 80.45 80.66 81.70 81.66
Effect of weight of Weak Noise Frozen Text feature.
We examine the sensitivity of ANPrompt to the weighting factor applied to the Weak Noise Frozen Text feature. As shown in Table 6, setting the weight to a small value (0.001) yields the best harmonic mean (81.70), indicating effective integration of weak noise signals. As the weight increases, performance on novel classes consistently degrades, leading to a gradual drop in HM. This trend suggests that overemphasizing the weak noise branch introduces distributional bias and harms generalization. These results highlight the necessity of a carefully balanced contribution from TextNoise.
Table 6: The Effectiveness of different weight of Weak Noise Frozen Text feature.
Weight Base Novel HM
0.001 86.15 77.70 81.70
0.01 86.10 76.84 81.21
0.1 86.09 76.48 81.00
1.0 86.02 75.51 80.43
Conclusion
We proposed ANPrompt, a robust prompt tuning framework that explicitly models weak semantic perturbations to improve generalization in vision-language models. By injecting weak semantic perturbation into the aligment process and prompt representation, ANPrompt enhances robustness under theweak semantic perturbation. Experimental results on 11 benchmarks demonstrate its clear advantages over existing methods, highlighting the effectiveness of incorporating weak semantic perturbation noise into prompt tuning.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62376106).

### v3

@@
-ANPrompt: Anti-noise Prompt Tuning for Vision-Language Models
+Robust Prompt Tuning for Vision-Language Models with Mild Semantic Noise
 Abstract
-Prompt tuning has emerged as an efficient and effective technique for adapting vision-language models (VLMs) with low computational overhead.
-However, existing methods often overlook the vulnerability of prompt-tuned VLMs to weak semantic perturbations—such as subtle image or text noise—that degrade their generalization to unseen classes.
-To address this limitation, we propose ANPrompt, a novel prompt tuning framework designed to enhance robustness under such perturbations.
-ANPrompt first constructs weak noise text features by fusing original and noise-perturbed text embeddings, which are then clustered to form noise prompts.
-These noise prompts are integrated with learnable prompt tokens to generate anti-noise prompts, which are injected into the deeper layers of both image and text encoders.
-To further capture the noise-aware visual semantics, ANPrompt computes the Noise-Resistant Visual Prompt Prototype (NRVPP) by averaging the output prompt tokens from the vision encoder.
-Finally, ANPrompt introduces alignment, robustness, and anti-noise objectives by computing a Weak semantic noise Alignment Loss (WALoss) alongside the standard cross-entropy and sim loss.
-Experiments across 11 benchmarks demonstrate that ANPrompt consistently outperforms existing prompt tuning approaches, achieving superior robustness to semantic noise and improved generalization to novel categories.
-Introduction
-Vision-language models (VLMs) like CLIP have demonstrated impressive performance on open-vocabulary recognition tasks  (Hua et al.
-2025; Weng et al.
-2023; Xu et al.
-2024), but adapting such large models to downstream tasks via full fine-tuning can be computationally prohibitive.
-To address this, prompt tuning (Zhu et al.
-2024; Lafon et al.
-2024; Shi et al.
-2024) has emerged as a lightweight alternative by optimizing only a small set of learnable tokens, achieving state-of-the-art results across various domains.
-Despite its efficiency, early prompt tuning methods such as CoOp (Zhou et al.
-2022b) often overfit to seen classes, resulting in limited generalization to novel categories and vulnerability to subtle semantic perturbations.
-To enhance transferability, recent approaches have integrated inductive biases into the prompting process.
-For instance, CoCoOp conditions prompts on image features to mitigate overfitting; KgCoOp constrains semantic drift by leveraging frozen text embeddings; and MaPLe adopts multimodal prompts to achieve more robust alignment.
-More advanced techniques such as PromptSRC and CoPrompt regularize prompt learning using frozen zero-shot logits, while MMRL and TAC construct shared or clustered semantic spaces.
-Although designed to reduce overfitting, consistency constraints based on frozen features may introduce noise information such as misleading text or visual clutter caused by domain gaps, ultimately weakening the robustness of Vision-Language Models.
-Figure 1: Weak semantic perturbations in image and text: CAMs may highlight irrelevant regions (e.g., background or co-occurring objects), while subtle textual cues (e.g., “dark stripes”) may mislead class semantics (e.g., shift from “cat” to “tiger”).
-Figure 2: Different with existing methods use frozen logits as constraint, risking overfitting under weak semantic perturbations.
-ANPrompt introduces weak semantic perturbations to generate soft logits and anti-prompt to improves generalization.
-As shown in Figure 1, we define subtle noise in images and texts as weak semantic perturbations, which appear as either misleading text (e.g., “a cat resembling a small tiger”) or visual noise (e.g., background clutter or co-occurring objects).
-While high-quality image-text pairs exhibit clear object focus and accurate descriptions, existing prompt tuning methods often struggle to handle such subtle shifts, leading to significant performance drops when adapting to novel categories.
-This highlights the urgent need for VLMs capable of learning robust, noise-resilient cross-modal representations to maintain alignment consistency and improve generalization.
-To address the issue of weak semantic perturbations, we propose ANPrompt — a prompt tuning framework specifically designed to enhance robustness under such perturbations.
-As illustrated in Figure 2, existing methods rely on frozen feature constraints and thus tend to suffer from overfitting when exposed to weak semantic perturbations.
-In contrast, our ANPrompt framework intentionally incorporates weak semantic perturbation information into prompt tokens and logits, thereby boosting the robustness of both the learned prompts and class distributions in scenarios involving weak semantic perturbations.
-Specifically, ANPrompt randomly samples two intra-class descriptions from a language cache: a main text that conveys core semantics, and a noise text that introduces minor linguistic variations.
-Their fusion generates a Weak Noise Frozen Text Feature, which incorporates weak semantic perturbation information while preserving class identity.
-These features are clustered into noise prompts and injected as biases into learnable tokens, forming Anti-noise Prompts.
-These prompts are then inserted into the higher layers of both text and vision encoders to introduce weak semantic perturbation information at the prompt representation level.
-After encoding, ANPrompt extracts image Feature, text Feature, and the Noise-Resistant Visual Prompt Prototype (NRVPP).
-The latter is obtained by mean-pooling the visual prompt tokens from the final layer of the visual encoder.
-Since Anti-noise Prompts are injected into deeper encoder layers, the resulting NRVPP inherit weak semantic perturbations.
-To stabilize training under both weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss (WALoss) between Weak Noise Logits and Logits set, which enforces alignment consistency under controlled weak semantic perturbations.
-Additionally, we employ a CE and sim loss to enable the model to learn task-relevant knowledge under weak semantic perturbations.
-By combining these losses, ANPrompt can improve the robustness of VLMs under weak semantic perturbations.
-Extensive experiments conducted on 11 benchmarks and base-to-new splits show that ANPrompt significantly outperforms existing prompt tuning methods, achieving state-of-the-art performance in terms of robustness and generalization.
+Prompt tuning has shown promising results, but its robustness and generalization to unseen categories remain limited.
+Through our experiments, we demonstrate that the complete removal of semantic noise is a key factor restricting robustness.
+Existing methods typically suppress or filter out semantic noise in the prompt space, inadvertently hindering the model’s robustness and its ability to generalize to unseen categories.
+To address this, we propose ANPrompt, a robust prompt tuning framework that actively incorporates weak semantic noise.
+By clustering weakly perturbed features into noise prompts and integrating them with learnable tokens in both the text and vision encoders, ANPrompt ensures controlled exposure to semantic variations.
+To enhance the visual pathway, we introduce the Noise-Resistant Visual Prompt Prototype (NRVPP), which stabilizes visual semantics under weak perturbations.
+Additionally, we propose a Weak Alignment Loss (WALoss) at the logits level to enforce consistency between clean and perturbed predictions, providing stable supervision.
+By combining weak semantic noise exposure with logits-based consistency, ANPrompt prevents overfitting to specific phrasings while preserving semantic integrity.
+Extensive experiments across 11 benchmarks, including base-to-new splits, show that ANPrompt consistently outperforms existing prompt tuning methods, offering superior robustness to semantic noise and improved generalization across tasks.
+Figure 1: This figure compares the accuracy of various methods, including DAPT, Synreplace, Mask, Shuffle, and Drop.
+The results show that incorporating weak semantic noise achieves superior performance, outperforming DAPT (79.35%) and other baseline methods.
+This highlights the effectiveness of weak semantic noise injection in enhancing model robustness and generalization.
+Figure 2: Framework comparison with representative filtering-based methods such as DAPT and ArGue.
+While these approaches constrain frozen logits and risk overfitting under weak semantic perturbations, the proposed ANPrompt leverages weak semantic perturbations to generate soft logits and anti-noise prompts, thereby improving generalization.
+Figure 3: Illustration of weak semantic perturbations in both image and text.
+Class activation maps (CAMs) may highlight irrelevant regions such as background or co-occurring objects, while subtle textual cues (e.g., “dark stripes”) can shift class semantics from “cat” to “tiger”.
+These examples demonstrate how weak semantic noise can mislead recognition and motivate the need for robust prompt learning.
+Figure 4: Quantitative comparison of different noise strategies in terms of Text Shift (TS), Logit Preservation Rate (LPR), and Accuracy Shift (AS).
+The results show that strong perturbations (drop, mask, shuffle, synonym replacement) cause semantic distortion and performance degradation, whereas our weak semantic perturbations (WSP) maintain negligible shift, high prediction consistency, and nearly zero accuracy loss.
+I Introduction
+Vision-language models (VLMs), such as CLIP, have demonstrated remarkable performance on open-vocabulary recognition tasks openworldauc ; openvclip ; xu2024llava .
+However, adapting these large-scale models to downstream tasks is often infeasible due to the high computational cost of full fine-tuning.
+Prompt tuning awt ; gallop ; qnet has emerged as an efficient alternative, updating only a small set of learnable tokens.
+Early approaches, such as CoOp coop , tend to overfit to seen classes, leading to poor generalization to novel categories and sensitivity to semantic perturbations.
+Subsequent methods, including CoCoOp cocoop , KgCoOp kgcoop , MaPLe maple , and MMRL mmrl , introduce inductive biases or shared semantic spaces to enhance robustness, but they still struggle under noisy conditions.
+Existing methods for handling noisy information in prompt tuning typically follow a filtering philosophy, where noise is suppressed or eliminated.
+Specifically, methods like DAPT dapt and ArGueargue align foreground features with textual descriptions while suppressing irrelevant background, as shown in Figure 2(a).
+However, these techniques often risk discarding valuable semantic variations and struggle to adapt to subtle, real-world perturbations.
+Furthermore, we demonstrate that filtering or suppressing noise, as done by methods like DAPT (79.35%) and Synreplace (81.44%), often limits robustness.
+These approaches suppress irrelevant background noise, which not only discards valuable semantic information but also hinders the model’s generalization and robustness to diverse real-world perturbations.
+In contrast, our method actively integrates weak semantic noise, resulting in a notable increase in performance.
+As shown in Figure 1, methods that filter or suppress noise show limited improvement, while our approach of incorporating weak semantic noise consistently outperforms these baseline methods.
+This emphasizes that actively incorporating weak semantic noise is more effective than filtering it, enabling the model to generalize better across unseen categories and enhancing its robustness to perturbations.
+As shown in Figure 2(b), our method actively injects weak semantic noise, enhancing robustness by integrating it with constraint-based techniques, rather than suppressing it.
+This allows the model to better adapt to diverse inputs without sacrificing semantic integrity.
+Noise injection has also been explored in the field of multimodal representation learning.
+Related works, such as noise1 , have proposed several text-based noise injection strategies, including strong perturbations like token dropping, masking, word shuffling, and synonym replacement, as illustrated in Figure 3.
+While these strong perturbations introduce variation, they often distort the original semantics and degrade model generalization, as demonstrated in the top-left part of the figure.
+However, we find that such strong noise injection methods are not well-suited for the Prompt tuning domain.
+To address this issue, we propose a more gradual noise injection approach, introducing weak semantic perturbations, which apply subtle shifts in both the image and text domains.
+By incorporating a logits constraint, shown in the bottom portion of Figure 3, our method ensures these perturbations preserve class identity while enhancing robustness.
+The stable Noise ZeroShot Logits generated by this approach help the model generalize better to unseen variations without the disruptive effects of strong noise perturbations.
+To substantiate this, we run a controlled comparison under the base-to-new setting, as shown in Figure 4.
+Strong perturbations severely distort semantics: token dropping yields $\mathrm{TS}=0.0196$ and $\mathrm{AS}=0.618$, masking further increases distortion with $\mathrm{TS}=0.0300$ and $\mathrm{AS}=0.923$, and word shuffling reaches $\mathrm{TS}=0.0573$ with $\mathrm{AS}=0.438$.
+Even synonym replacement introduces instability ($\mathrm{TS}=0.0090$, $\mathrm{AS}=0.00013$).
+In contrast, weak semantic noise preserves class identity, exhibiting negligible text shift ($\mathrm{TS}=6.3\times 10^{-6}$ at $\epsilon=0.01$), high prediction consistency ($\mathrm{LPR}\approx 0.98$), and nearly zero accuracy degradation.
+These results suggest that controlled weak noise exposure can enhance robustness without sacrificing semantic fidelity.
+Building on these insights, we propose ANPrompt, a prompt-tuning framework that actively leverages weak semantic noise rather than suppressing it.
+Specifically, ANPrompt clusters weak noise features into representative prompts and integrates them with learnable tokens in both text and vision encoders.
+As shown in Figure 2(b), our approach contrasts with previous methods that focus on noise suppression.
+By actively incorporating weak semantic noise, the model improves its robustness to unseen categories, while effectively enhancing generalization and maintaining semantic consistency.
+To further capture noise-aware visual semantics, we introduce a Noise-Resistant Visual Prompt Prototype (NRVPP), which provides stable visual representations under weak perturbations.
+This controlled noise exposure regularizes prompt representations, mitigates overfitting to specific textual expressions, and promotes learning of more robust class-level semantics.
+Furthermore, we introduce a Weak Alignment Loss (WALoss) that enforces consistency between clean and perturbed predictions.
+Unlike feature-level perturbations, which may disrupt intermediate representations and lead to instability, enforcing consistency at the logits level provides a more stable supervision signal.
+This directly regularizes the decision boundary and improves robustness without compromising the model’s discriminative capacity.
+The benefits of WALoss align with our overall approach of controlled noise exposure, further enhancing the model’s ability to adapt to real-world variations without losing accuracy or generalization.
 Our main contributions are summarized as follows:
-  • We propose ANPrompt, a prompt tuning framework that improve the robustness of VLMs by injecting weak semantic perturbations into both prompt tokens and logits pathways.
-  • We introduce a Weak semantic noise Alignment Loss (WALoss) to enforce alignment consistency under controlled weak semantic perturbations.
-  • We develop a weak semantic perturbations prompt injection mechanism that clusters weak noise frozen text features into noise prompts and injects them into learnable tokens to get anti-noise prompt, which is used to inject weak semantic perturbations at the prompt representation level.
-  • ANPrompt achieves state-of-the-art performance on 11 benchmark datasets under base-to-new generalization settings.
-Related Work
-Vision-Language Models
-Pretrained vision-language models (VLMs) such as CLIP (Radford et al.
-2021) and ALIGN (Jia et al.
-2021) achieve strong open-vocabulary and cross-modal capabilities through contrastive learning on large-scale image-text pairs.
-Building on this foundation, models like BLIP (Li et al.
-2022), FILIP (Yao et al.
-2021), Florence (Yuan et al.
-2021), and LiT (Zhai et al.
-2022) enhance fine-grained alignment and data efficiency via retrieval augmentation(Li et al.
-2024a), self-supervised learning, and stronger backbones.
-NLIP (Huang et al.
-2023) further improves robustness by introducing noise-harmonization based on alignment uncertainty and noise-completion via concept-guided caption generation.
+  • We propose ANPrompt, a prompt tuning framework that improves robustness by incorporating weak semantic noise into both prompt tokens and logits.
+  • We introduce a Weak Alignment Loss (WALoss) that enforces prediction consistency at the logits level, providing stable supervision and stronger generalization compared to feature-level perturbations.
+  • We exploit weak semantic noise as a source of natural linguistic diversity, preserving class identity while preventing overfitting to specific phrasings, thereby enhancing base-to-new generalization compared with strong noise strategies.
+  • We conduct extensive experiments on 11 datasets and base-to-new splits, demonstrating that ANPrompt achieves state-of-the-art robustness and generalization compared with existing prompt tuning methods.
+II Related Works
+II-A Vision-Language Models
+Pretrained vision-language models (VLMs)l1 ; l2 ; l3 ; l4 ; l5 such as CLIP clip and ALIGN align achieve strong open-vocabulary and cross-modal capabilities through contrastive learning on large-scale image-text pairs.
+Building on this foundation, models like BLIP blip , FILIP filip , Florence florence , and LiT lit enhance fine-grained alignment and data efficiency via retrieval augmentationli2024freestyleret , self-supervised learning, and stronger backbones.
+NLIP NLIP further improves robustness by introducing noise-harmonization based on alignment uncertainty and noise-completion via concept-guided caption generation.
 These advances enable effective zero-shot and few-shot transfer for tasks such as retrieval, classification, and grounding.
-However, adapting such large models to downstream tasks (Li et al.
-2025a, 2023) remains costly due to domain shifts and the high overhead of full fine-tuning.
-To address this, parameter-efficient tuning methods—such as adapter-based approaches (Gao et al.
-2024; Yang et al.
-2024b; Song et al.
-2023) and prompt learning (Zhou et al.
-2022b, a; Khattak et al.
-2024)—have been proposed to enable lightweight adaptation without retraining the entire model.
-Prompt Learning
+However, adapting such large models to downstream tasks li2025beyond ; li2023weakly ; pcl ; hipl ; sod remains costly due to domain shifts and the high overhead of full fine-tuning.
+To address this, parameter-efficient tuning methods—such as adapter-based approaches clipadapter ; mma ; metaadapter and prompt learning protext ; cpr ; pmp —have been proposed to enable lightweight adaptation without retraining the entire model.
+II-B Prompt Learning
 Prompt learning has become an efficient strategy to adapt VLMs without full fine-tuning.
-Early methods such as CoOp (Zhou et al.
-2022b) and CoCoOp (Zhou et al.
-2022a) explore soft and image-conditioned prompts to enhance generalization.
-Subsequent works improve robustness and transferability through multimodal regularization (MaPLe (Khattak et al.
-2023a), PromptSRC (Khattak et al.
-2023b)), gradient alignment (ProGrad (Zhu et al.
-2023)), hierarchical and knowledge-guided prompts (HPT (Wang et al.
-2024), TCP (Yao, Zhang, and Xu 2024)), and distillation-based initialization (CLIP-kd (Yang et al.
-2024a), ComKD-CLIP (Chen et al.
-2024), PromptKD (Li et al.
-2024b), CasPL (Wu et al.
-2024)).
-Other methods focus on attribution-based prompting (ATPrompt (Li et al.
-2024c)), two-stage decoupled optimization (DPC (Li et al.
-2025b), DePT (Zhang et al.
-2024), 2SFS (Farina et al.
-2025)), or cross-modal adapters (MMA (Yang et al.
-2024b)) for few-shot settings.
-Consistency learning (CoPrompt (Roy and Etemad 2024)) and task-aware clustering (TAC (Hao et al.
-2025)) further enhance adaptability.
+Early methods such as CoOp coop and CoCoOp cocoop explore soft and image-conditioned prompts to enhance generalization.
+Subsequent works improve robustness and transferability through multimodal regularization (MaPLe maple , PromptSRC promptsrc ), gradient alignment (ProGrad prograd ), hierarchical and knowledge-guided prompts (HPT hpt , TCP tcp ), and distillation-based initialization (CLIP-kd clipkd , ComKD-CLIP comkdclip , PromptKD promptkd , CasPL caspl ).
+Other methods focus on attribution-based prompting (ATPrompt atprompt ), two-stage decoupled optimization (DPC dpc , DePT dept , 2SFS 2sfs ), or cross-modal adapters (MMA mma ) for few-shot settings.
+Consistency learning (CoPrompt coprompt ) and task-aware clustering (TAC tac ) further enhance adaptability.
 However, these approaches largely target generalization and efficiency, with limited attention to robustness against weak semantic perturbations.
-Prompting under Weak Semantic Perturbation Noise
+II-C Prompting under Weak Semantic Perturbation Noise
 Recent works have explored robust prompt learning under noisy conditions.
-Methods such as NLPrompt (Pan et al.
-2025), JoAPR (Guo and Gu 2024), and ArGue (Tian et al.
-2024) mitigate label noise through loss design, sample filtering, and LLM-generated attributes, while ClipCleaner (Feng, Tzimiropoulos, and Patras 2024) leverages CLIP-based sample selection to decouple training from noise filtering.
+Methods such as NLPrompt nlprompt , JoAPR joapr , and ArGue argue mitigate label noise through loss design, sample filtering, and LLM-generated attributes, while ClipCleaner clipcleaner leverages CLIP-based sample selection to decouple training from noise filtering.
 However, these approaches mainly focus on filtering noise, which risks discarding informative variations and often limits generalization.
@@
 By integrating perturbed prompts with clustering-based bias representations, ANPrompt enables robust and discriminative prompt learning without relying on brittle filtering heuristics.
-Figure 3: ANPrompt randomly samples a main and a noise sentence from the same class to construct the Weak Noise Frozen Text Feature, which is used to generate Anti-noise Prompts.
-These prompts are injected into the encoder’s deep layers.
-During encoding, we extract the Prompted Image/Text Features and the Noise-Resistant Visual Prompt Prototype (NRVPP) to compute four types of logits, which are supervised by WALoss and auxiliary losses to improve robustness against weak semantic perturbations.
-Methodology
-Preliminary
-Vision-language models (VLMs) such as CLIP (Radford et al.
-2021) learn a joint embedding space for images and texts through contrastive learning over large-scale image-text pairs.
+III Method
+III-A Preliminaries
+Vision-language models (VLMs) such as CLIP clip learn a joint embedding space for images and texts through contrastive learning over large-scale image-text pairs.
 These models exhibit strong zero-shot generalization but are highly sensitive to semantic variations in either modality.
 A typical VLM consists of a vision encoder and a text encoder.
-The vision encoder tokenizes the input image into patches, encodes them through Transformer layers, and outputs a class token representing global semantics.
+The vision encoder tokenizes an input image into patches, encodes them through Transformer layers, and outputs a class token representing global semantics.
 The text encoder processes class-dependent prompts (e.g., “a photo of a [CLASS]”), encodes them with Transformers, and uses the [EOS] token as the prompt embedding.
@@
 where $f$ is the image feature, $w_{c}$ is the text feature for class $c$, $\tau$ is a learnable temperature, and $\text{sim}(\cdot)$ denotes cosine similarity.
-Overview
-We propose ANPrompt — a robust prompt tuning framework that injects weak semantic perturbations into the prompting and alignment processes to enhance the robustness of VLMs under such perturbations.
-Specifically, ANPrompt simulates weak semantic perturbations by fusing two intra-class textual descriptions from the language cache, generating a Weak Noise Frozen Text Feature.
-These Weak Noise Frozen Text Features are clustered to obtain noise prompts, which are then projected and added to learnable prompt tokens to form Anti-noise Prompts.
-These prompts are inserted into the deeper layers of both vision and text encoders, enabling ANPrompt to inject weak semantic perturbations at the prompt level.
-Subsequently, prompted image features, prompted text features, and NRVPP can be obtained for computing alignment logits, robustness logits, weak noise logits, and anti-noise logits.
-During training, ANPrompt combines the Weak Semantic Noise Alignment Loss (WALoss), cross-entropy loss, and SimLoss to enhance the robustness of VLMs under weak semantic perturbations.
-ANPrompt Framework
-As shown in Figure 3.
-The ANPrompt framework proceeds through four sequential stages: (1) Weak Noise Frozen Text Feature Construction, (2)Anti-noise Prompt Generation, (3) Feature Encoding, and (4) Logit Computation.
-1.
-Weak Noise Frozen Text Feature Generation: ANPrompt samples two intra-class descriptions: a main text that encodes core semantics, and a noise text that introduces subtle phrasing variations (e.g., attribute reordering).
-These two texts are encoded by the frozen CLIP text encoder into $f_{m}$ and $f_{n}$ respectively, and fused as follows:
-$f_{w}=f_{m}+\alpha f_{n}$ (2)
-where $\alpha$ regulates the perturbation strength.
-Since the Weak Noise Frozen Text Feature $f_{w}$ incorporates contextual information from different sentences, it inevitably captures contextual cues that lean toward semantically adjacent yet incorrect classes.
-Consequently, weak noise text features inherently encode weak semantic perturbations information.
-2.
-Anti-noise Prompt Generation: To map weak semantic perturbation information into the prompt space, ANPrompt performs K-means clustering on all $f_{w}$ to extract representative semantic centers, which are referred to as noise prompts.
-These noise prompts are added to the learnable prompt tokens $P_{c}$ to form intermediate anti-noise prompts $P_{a}$.
-These intermediate prompts are then fed into two separate fully connected layers to generate modality-specific Anti-noise Prompts, which is injected into the deeper layers of both encoders.
-3.
-Feature Encoding.
-After prompt injection, we extract three types of features: the Prompted Image Feature $f_{v}$, encoded from the final image class token; the Prompted Text Feature $f_{t}$, derived from the final [EOS] token; and the Noise-Resistant Visual Prompt Prototype (NRVPP) $f_{N}$, obtained by mean-pooling the visual Anti-noise Prompts after encoding, which captures weak semantic perturbations information.
-4.
-Logit Computation.
-We compute four sets of logits: the Alignment Logits $\ell_{A}=f_{v}\cdot f_{t}^{\top}$, which measure the similarity between the prompted image and text features; the Robustness Logits $\ell_{R}=f_{v}\cdot f_{w}^{\top}$, which assess the model’s robustness against weak semantic noise in text; the Weak Noise Logits $\ell_{W}=f_{N}\cdot f_{w}^{\top}$, which reflect the alignment between noise-resistant visual features and weakly perturbed text features; and the Anti-noise Logits $\ell_{\text{final}}=\theta\ell_{A}+(1-\theta)\ell_{R}$, which integrate alignment and robustness for final prediction.
-To stabilize training under weak semantic perturbations, we apply $\mathcal{L}_{\text{WA}}$ between $\ell_{W}$ and $\ell_{A}$ for alignment consistency under weak semantic perturbations, $\mathcal{L}_{\text{CE}}$ on $\ell_{\text{final}}$ for task supervision, and $\mathcal{L}_{\text{sim}}$ between frozen and prompted features to enhance semantic consistency.
-The final loss is:
-$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{sim}}+\gamma\mathcal{L}_{\text{WA}}$ (3)
-where $\mathcal{L}_{\text{CE}}$ denotes a cross-entropy loss, $\mathcal{L}_{\text{sim}}$ represents a cosine loss, and $\mathcal{L}_{\text{WA}}$ is defined as follows.
-$\lambda$ serves as the weight of $\mathcal{L}_{\text{sim}}$, while $\gamma$ is a variance-adaptive weight used to regulate the strength of weak semantic perturbations.
-Weak Noise Alignment Loss
-To enforce alignment consistency under weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss ($\mathcal{L}_{\text{WA}}$) that minimizes the divergence between $\ell_{W}$ and $\ell_{A}$.
-Since KL divergence measures the discrepancy between distributions, minimizing it encourages the perturbed prediction $\ell_{W}$ to stay close to the original alignment $\ell_{A}$, penalizing semantic shifts and improving robustness while preserving alignment structure.
-To compute $\mathcal{L}_{\text{WA}}$, we first derive $\ell_{A}$ from the prompted image and text features.
-Then, to obtain $\ell_{W}$, we compute the noise-resistant visual prototype $f_{N}$ by averaging the visual prompt token features $f_{v}^{o}\in\mathbb{R}^{N\times C}$:
-$f_{N}=\frac{1}{N}\sum_{i=1}^{N}f_{v}^{o(i)}$ (4)
-This averaging suppresses token-level variations and mitigates visual semantic distraction, especially under anti-prompts with weak semantic noise, where individual tokens may attend to irrelevant regions such as backgrounds or co-occurring objects.
-The weak noise logits are then computed as:
-$\ell_{W}=f_{N}\cdot f_{w}^{\top}$ (5)
-Finally, the weak alignment loss is defined as:
-$\mathcal{L}_{\text{WA}}=\gamma\cdot\text{KL}(\text{softmax}(\ell_{R})\|\text{softmax}(\ell_{W}))$ (6)
-with the variance-adaptive weight $\gamma$ given by:
-$\gamma=\frac{1}{\text{std}(\ell_{R})\cdot|\ell_{R}|+\epsilon_{0}}$ (7)
-where $\text{std}(\ell_{R})$ denotes the standard deviation over all elements in $\ell_{R}$, $|\ell_{R}|$ represents the total number of elements (i.e., numel() of logits), and $\epsilon_{0}$ is a small constant to prevent division by zero.
-This global scaling dynamically adjusts the loss weight based on the dispersion and size of the logits, stabilizing optimization under varied prediction confidence levels.
-Weak Semantic Perturbations Prompt Injection Mechanism
-We introduce a prompt injection mechanism based on weak semantic perturbations.
-Specifically, we first cluster the weak noise frozen text features $f_{w}\in\mathbb{R}^{N\times C}$ into $K$ noise prompts $P_{w}\in\mathbb{R}^{k\times C}$ using K-Means, where $k$ matches the number of learnable prompt tokens $P_{c}\in\mathbb{R}^{k\times C}$ to ensure one-to-one integration.
-These noise prompts are added to the learnable tokens to form the anti-noise prompt:
-$P_{a}=P_{c}+\epsilon P_{w}$ (8)
-where $\epsilon$ controls the strength of injected weak perturbations. The anti-noise prompt $P_{a}$ is then projected into modality-specific spaces via a mapping layer.
-Finally, the projected prompts $P_{a}^{\text{image}}$ and $P_{a}^{\text{text}}$ are injected into the visual and textual input sequences:
-$\{P_{a}^{\text{image}},e_{\text{cls}},e_{1},e_{2},\dots,e_{M}\}$ (9)
-$\{t_{\text{SOS}},P_{a}^{\text{text}},t_{1},t_{2},\dots,t_{L},c_{k},t_{\text{EOS}}\}$ (10)
-This injection enriches the prompt space with weakly perturbed semantic cues, promoting robustness against subtle noise while maintaining structural alignment with the original prompt format.
-Table 1: Comparison with state-of-the-art methods on base-to-novel generalization.
+Figure 5: Overview of ANPrompt.
+For each class, a main and a noise sentence are sampled to construct the Weak Noise Frozen Text Feature, which is clustered into noise prompts.
+These are combined with learnable tokens to generate Anti-noise Prompts, which are injected into the last two layers of both vision and text encoders.
+During encoding, we obtain Prompted Image/Text Features and the Noise-Resistant Visual Prompt Prototype (NRVPP) to compute four types of logits, supervised by WALoss and auxiliary losses to improve robustness against weak semantic perturbations.
+III-B ANPrompt Framework
+We propose ANPrompt, a robust prompt tuning framework that explicitly leverages weak semantic noise to enhance the robustness of VLMs.
+Specifically, ANPrompt simulates weak perturbations by fusing intra-class textual descriptions, generating Weak Noise Frozen Text Features.
+These features are clustered to form noise prompts, which are added to learnable prompt tokens to construct Anti-noise Prompts.
+The Anti-noise Prompts are injected into the last two transformer blocks of both vision and text encoders by prepending them as additional tokens, thereby exposing the model to weak semantic perturbations during prompt learning.
+After encoding, we obtain prompted image features, prompted text features, and the Noise-Resistant Visual Prompt Prototype (NRVPP), which are used to compute four types of logits: alignment, robustness, weak noise, and anti-noise logits.
+During training, ANPrompt combines the Weak Alignment Loss (WALoss) with cross-entropy and similarity losses to enhance robustness against weak semantic perturbations.
+Specifically, WALoss encourages the model to maintain prediction consistency between clean and perturbed inputs at the logits level, ensuring that the model remains stable even when exposed to subtle changes in the data.
+By incorporating weak semantic noise, ANPrompt regularizes the decision boundary, making it more resilient to small semantic shifts without sacrificing accuracy.
+Unlike previous methods, which primarily employ generic prompt regularization based on cosine similarity or feature-level perturbations, ANPrompt takes a more explicit approach by integrating weak semantic noise directly into prompt tokens.
+It enforces consistency at the logits level through WALoss, which provides a more stable supervision signal compared to traditional methods.
+This mechanism is distinctive in its ability to simultaneously preserve class identity and improve generalization, making ANPrompt a more effective solution for handling noise in vision-language models.
+As illustrated in Figure 5, the ANPrompt framework proceeds through four sequential stages: (1) Weak Noise Frozen Text Feature Generation, (2) Anti-noise Prompt Construction, (3) Feature Encoding, and (4) Logit Computation.
+III-B1 Weak Noise Frozen Text Feature Generation
+ANPrompt samples two intra-class descriptions: a main text that encodes the core semantics and a noise text that introduces subtle phrasing variations (e.g., attribute reordering).
+These two texts are encoded by the frozen CLIP text encoder into features $f_{m}$ and $f_{n}$, which are then fused as:
+$f_{w}=f_{m}+\alpha f_{n},$ (2)
+where $\alpha$ is a scaling factor that controls the strength of the perturbation.
+The Weak Noise Frozen Text Feature $f_{w}$ combines contextual information from both the main and noise texts, capturing weak semantic perturbations that subtly shift semantics toward adjacent, but incorrect, classes.
+The main and noise texts are randomly sampled from a diverse intra-class language cache.
+This approach effectively captures natural linguistic variations, approximating real-world weak perturbations.
+Future work may explore more controlled alternatives, such as paraphrasing or synonym replacement, to generate more fine-grained noise variations.
+III-B2 Anti-noise Prompt Construction
+To map weak semantic perturbation information into the prompt space, ANPrompt applies K-means clustering on all $f_{w}$ to extract representative semantic centers, denoted as noise prompts.
+These noise prompts are then added to the learnable prompt tokens $P_{c}$ to form intermediate anti-noise prompts $P_{a}$.
+Two modality-specific fully connected layers then transform $P_{a}$ into Anti-noise Prompts for the vision and text encoders, which are injected into their deeper layers.
+III-B3 Feature Encoding
+After prompt injection, three types of features are obtained:
+  • Prompted Image Feature $f_{v}$, encoded from the final image class token;
+  • Prompted Text Feature $f_{t}$, derived from the [EOS] token;
+  • Noise-Resistant Visual Prompt Prototype (NRVPP) $f_{N}$, obtained by mean-pooling the visual Anti-noise Prompts.
+This step helps capture weak semantic perturbations while suppressing high-frequency token-level variations, ensuring the preservation of meaningful visual information across perturbations.
+III-B4 Logit Computation
+We compute four sets of logits:
+  • Alignment Logits $\ell_{A}=f_{v}\cdot f_{t}^{\top}$, measuring the similarity between prompted image and text features;
+  • Robustness Logits $\ell_{R}=f_{v}\cdot f_{w}^{\top}$, assessing the model’s robustness to weak semantic noise in the text;
+  • Weak Noise Logits $\ell_{W}=f_{N}\cdot f_{w}^{\top}$, reflecting the alignment between noise-resistant visual features and weakly perturbed text features;
+  • Anti-noise Logits $\ell_{\text{final}}=\theta\ell_{A}+(1-\theta)\ell_{R}$, integrating alignment and robustness for the final prediction.
+After computing these four logits, we can compute the training losses: cross-entropy loss ($\mathcal{L}_{\text{CE}}$), weak alignment loss ($\mathcal{L}_{\text{WA}}$), and cosine similarity loss ($\mathcal{L}_{\text{sim}}$).
+Specifically:
+  • The cross-entropy loss ($\mathcal{L}_{\text{CE}}$) is computed using the Alignment Logits $\ell_{A}$.
+  • The weak alignment loss ($\mathcal{L}_{\text{WA}}$) is computed based on the Weak Noise Logits $\ell_{W}$ and the Alignment Logits $\ell_{A}$.
+  • The cosine similarity loss ($\mathcal{L}_{\text{sim}}$) is calculated between the image and text embeddings to preserve semantic consistency across modalities.
+The overall training loss is:
+$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{sim}}+\gamma\mathcal{L}_{\text{WA}},$ (3)
+where $\lambda$ controls the contribution of $\mathcal{L}_{\text{sim}}$, and $\gamma$ adaptively controls the effect of $\mathcal{L}_{\text{WA}}$.
+where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{sim}}$ is the cosine similarity loss, and $\mathcal{L}_{\text{WA}}$ is the weak alignment loss.
+Here, $\mathcal{L}_{\text{sim}}$ is defined over both image and text embeddings to preserve semantic consistency across modalities.
+In addition, $K$ explicitly denotes the number of learnable prompt tokens, ensuring one-to-one integration with the clustered noise prompts.
+III-C Weak Alignment Loss
+To enforce consistency under weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss ($\mathcal{L}_{\text{WA}}$) that minimizes the divergence between $\ell_{W}$ and $\ell_{A}$.
+Minimizing this divergence encourages predictions under perturbation to remain close to the clean alignment, penalizing semantic shifts while preserving structural consistency.
+To compute $\ell_{W}$, the NRVPP $f_{N}$ is obtained by averaging visual prompt token features $f_{v}^{o}\in\mathbb{R}^{N\times C}$:
+$f_{N}=\frac{1}{N}\sum_{i=1}^{N}f_{v}^{o(i)},$ (4)
+which suppresses token-level variations and mitigates visual semantic distraction caused by weak noise. The weak noise logits are then:
+$\ell_{W}=f_{N}\cdot f_{w}^{\top}.$ (5)
+The Weak Alignment Loss is defined as:
+$\mathcal{L}_{\text{WA}}=\text{KL}(\text{softmax}(\ell_{A})\,\|\,\text{softmax}(\ell_{W})),$ (6)
+We use $\gamma$ to control the variance-adaptive weight:
+$\gamma=\frac{1}{\text{std}(\ell_{R})\cdot|\ell_{R}|+\epsilon_{0}},$ (7)
+where $\text{std}(\ell_{R})$ is the standard deviation of $\ell_{R}$, $|\ell_{R}|$ is its number of elements, and $\epsilon_{0}$ is a small constant for numerical stability.
+This dynamic scaling adjusts the loss weight based on prediction dispersion, stabilizing optimization across varied confidence levels.
+III-D Weak Semantic Perturbation Prompt Injection
+Finally, we formalize the injection mechFanism.
+The weak noise features $f_{w}\in\mathbb{R}^{N\times C}$ are clustered into $K$ noise prompts $P_{w}\in\mathbb{R}^{K\times C}$ using K-means, with $K$ matching the number of learnable prompt tokens $P_{c}\in\mathbb{R}^{K\times C}$.
+The anti-noise prompts are then defined as:
+$P_{a}=P_{c}+\epsilon P_{w},$ (8)
+where $\epsilon$ controls the perturbation strength.
+$P_{a}$ is projected into modality-specific spaces to obtain $P_{a}^{\text{image}}$ and $P_{a}^{\text{text}}$, which are injected into the input sequences:
+$\{P_{a}^{\text{image}},e_{\text{cls}},e_{1},\dots,e_{M}\},$ (9)
+$\{t_{\text{SOS}},P_{a}^{\text{text}},t_{1},\dots,t_{L},c_{k},t_{\text{EOS}}\}.$
+This injection enriches the prompt space with weakly perturbed cues, promoting robustness against subtle noise while maintaining structural alignment with the original prompt format.
+TABLE I: Comparison with state-of-the-art methods on base-to-novel generalization.
 The best accuracies are bold.
 We can see that our method achieves the best performance compared with other state-of-the-art methods.
-Method
-Ref.
-Average ImageNet Caltech101 OxfordPets
+Method Ref. Average ImageNet Caltech101 OxfordPets
 Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
 MMRL CVPR25 85.68 77.16 81.20 77.90 71.30 74.45 98.97 94.50 96.68 95.90 97.60 96.74
+SurPL-G ICML25 86.37 76.32 81.03 78.74 70.49 74.39 98.77 95.16 96.93 96.37 97.41 96.89
+LwEIB IJCV25 84.45 78.21 81.21 76.64 71.64 74.06 98.47 95.47 96.95 95.70 97.40 96.54
+BIP-D TPAMI25 85.95 75.79 80.55 78.27 69.90 73.85 98.77 94.47 96.57 95.30 97.37 96.32
 ANPrompt 86.15 77.70 81.70 77.83 71.17 74.35 98.97 94.73 96.80 95.73 97.17 96.44
-Method
-Ref.
-StandfordCars Flowers102 Food101 FGVCAircraft
+Method Ref. StandfordCars Flowers102 Food101 FGVCAircraft
 Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
 MMRL CVPR25 81.30 75.07 78.06 98.97 77.27 86.78 90.57 91.50 91.03 46.30 37.03 41.15
+SurPL-G ICML25 83.57 72.77 77.80 98.90 72.88 83.92 90.92 91.81 91.36 49.20 36.93 42.19
+LwEIB IJCV25 80.07 74.01 76.92 97.53 77.50 86.37 90.63 91.73 91.18 45.11 42.60 43.82
+BIP-D TPAMI25 83.27 73.00 77.80 98.30 75.90 85.66 90.73 91.23 90.98 48.83 35.90 41.38
 ANPrompt 83.57 74.63 78.85 98.60 77.30 86.66 90.63 91.50 91.06 49.67 36.60 42.14
-Method
-Ref.
-SUN397 DTD EuroSAT UCF101
+Method Ref. SUN397 DTD EuroSAT UCF101
 Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
 MMRL CVPR25 83.20 79.30 81.20 85.67 65.00 73.82 95.60 80.17 87.21 88.10 80.07 83.39
+SurPL-G ICML25 83.43 78.96 81.13 86.07 62.04 72.11 94.63 81.33 87.48 89.44 79.74 84.31
+LwEIB IJCV25 81.10 79.80 80.44 82.87 67.83 74.60 95.00 80.01 86.86 85.73 82.37 84.02
+BIP-D TPAMI25 83.07 79.13 81.05 85.43 58.77 69.63 95.70 78.00 85.95 87.77 80.00 83.70
 ANPrompt 83.07 79.07 81.02 85.20 65.10 73.80 95.53 87.33 91.21 88.80 80.07 84.21
-Table 2: Results from cross-dataset evaluation. ANprompt provides the highest average accuracy.
+TABLE II: Results from cross-dataset evaluation. ANPrompt provides the highest average accuracy.
 Source Target
@@
 TAC 72.77 94.53 90.67 65.30 72.20 85.83 23.53 67.63 47.57 48.07 70.00 66.53
+SurPL-G 73.33 93.73 90.16 64.67 70.53 85.52 24.80 67.43 48.54 52.85 67.86 66.61
+BIP 70.83 93.93 90.13 65.53 71.43 86.27 26.47 67.23 46.00 48.90 69.83 66.57
 ANPrompt 71.13 94.70 91.00 66.00 72.87 86.23 26.10 67.47 45.83 51.97 69.27 67.14
-Experiments
-Evaluation Task
-Base-to-Novel Generalization
-To assess the zero-shot generalization capability of ANprompt, we divide each dataset into base and novel categories.
-The model is trained on the base categories using a few-shot setting, and its performance is evaluated on both the base and novel categories without additional training.
-Cross-Dataset Evaluation
-To examine the transferability of the model across datasets, we conduct experiments where the model is trained on a few-shot subset of the ImageNet dataset and directly evaluated on other datasets, without any further adaptation or fine-tuning.
-Domain Generalization
-To test the model’s resilience to domain shift, we use the ImageNet-trained model and evaluate it on four distinct ImageNet-derived datasets that introduce various domain differences, allowing us to measure performance under out-of-distribution scenarios.
-Evaluation Settings
-Datasets
-In the tasks of Base-to-New Generalization and Cross-Dataset Evaluation, we followed the experimental setup of CoOp (Zhou et al.
-2022b), and evaluated on 11 public datasets, including ImageNet, Caltech (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al.
-2012), Flowers (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Gool 2014), StanfordCars (Krause et al.
-2013), FGVCAircraft (Maji et al.
-2013), EuroSAT (Helber et al.
-2019), UCF101 (Kay et al.
-2017), DTD (Cimpoi et al.
-2014) and SUN397 (Xiao et al.
-2010).
-We also use ImageNet (Deng et al.
-2009) as the source dataset, and use its four variants ImageNetV2 (Recht et al.
-2019), ImageNet-Sketch (Wang et al.
-2019), ImageNet-A (Hendrycks et al.
-2021b) and ImageNet-R (Hendrycks et al.
-2021a) as the target dataset to evaluate the domain generalization task.
-Implementation Details
+IV EXPERIMENT
+IV-A Dataset
+In the tasks of Base-to-New Generalization and Cross-Dataset Evaluation, we followed the experimental setup of CoOp coop , and evaluated on 11 public datasets, including ImageNet, Caltech caltech101 , OxfordPets oxfordpets , Flowers Flowers102 , Food101 food101 , StanfordCars StanfordCars , FGVCAircraft fgvcaircraft , EuroSAT Eurosat , UCF101 ucf101 , DTD dtd and SUN397 sun .
+We also use ImageNet imagenet as the source dataset, and use its four variants ImageNetV2 imagenetv2 , ImageNet-Sketch imagenets , ImageNet-A imageneta and ImageNet-R imagenetr as the target dataset to evaluate the domain generalization task.
+IV-B Implementation Details
 We set the learning rate to 0.001 and adopt the Adam optimizer to train ANPrompt.
 The model is trained for 10 epochs with a batch size of 4, using ViT-B/16 as the vision backbone.
-For the LLM cache, we incorporate prompts from CoPrompt (Roy and Etemad 2024) and HPT (Wang et al.
-2024).
+For the LLM cache, we incorporate prompts from CoPrompt coprompt and HPT hpt .
 And we set the $\theta$ and $\alpha$ with 0.7 and 0.001.
+We inject the prompt into the layer 6 to 12.
 All experiments are conducted on a single NVIDIA RTX 3090 Ti GPU.
 Reported results are averaged over three random seeds.
-Experimental Performance
-We compare the results of our method on three evaluation tasks with other advanced prompt learning methods, including zero-shot CLIP and some prompt learning methods, such as CoOp (Zhou et al.
-2022b), CoCoOp (Zhou et al.
-2022a), MaPLe (Khattak et al.
-2023a), PromptSRC (Khattak et al.
-2023b), CoPrompt (Roy and Etemad 2024), TextRefiner (Xie et al.
-2025) and MMRL (Guo and Gu 2025).
-We also added some targeted ablation experiments to prove the effectiveness of the proposed ANPrompt method.
-Base-to-New Generalization
-Table 1 shows the performance of ANPrompt and other comparison methods on B2N generalization tasks.
-Compared with the most advanced MMRL method, ANPrompt improves the base class by 0.44% and the new class by 0.43%.
-Combining the base class and the new class, ANPrompt achieves a performance improvement of 0.44% on the harmonic average.
-It is worth noting that compared with MMRL (Guo and Gu 2025), ANPrompt achieves significant performance improvement on EuroSAT dataset: 6.03% on new class and 3.31% on harmonic average.
+IV-C Base-to-New Generalization
+In this experiment, we evaluate ANPrompt and benchmark its performance against several existing methods, such as CLIP clip , CoOp coop , CoCoOp cocoop , MaPLe maple , PromptSRC promptsrc , CoPrompt coprompt , TextRefiner textrefiner , MMRL mmrl , SurPL-G surpl , LwEIB lweib and BIP-D bip .
+TABLE I shows the performance of ANPrompt and other comparison methods on B2N generalization tasks.
+Compared with the most advanced LwEIB method, ANPrompt improves the base class by 1.7% and the harmonic average by 0.49%.
+It is worth noting that compared with BIP-D, ANPrompt achieves significant performance improvement on EuroSAT dataset: 9.33% on new class and 5.26% on harmonic average.
 By transforming weak semantic noise into constructive training signals, ANPrompt achieves performance improvement by fusing complementary semantic perturbation sentences.
-Cross-Dataset Evaluation
-Table 2 shows the performance comparison between our ANPrompt and existing methods in cross-dataset evaluation.
+IV-D Cross-Dataset Generalization
+TABLE II shows the performance comparison between our ANPrompt and existing methods in cross-dataset evaluation.
 On the ImageNet source dataset, ANPrompt shows comparable performance to other competitive methods.
-On the target domain dataset, ANPrompt achieved the best performance on six of the ten datasets, showing excellent generalization ability, especially on fine-grained datasets, such as FGVCAircraft and EuroSAT.
+On the target domain dataset, ANPrompt achieved the best performance on four of the ten datasets, showing excellent generalization ability, especially on fine-grained datasets, such as FGVCAircraft and EuroSAT.
 Specifically, with an average of 10 datasets, ANPrompt achieved the highest accuracy rate of 67.14%.
@@
 Different from the existing methods, we use the anti-noise visual prompt prototype to align the weak noise frozen text features to generate soft, context-aware supervision signals, thus improving cross-domain performance.
-Table 3: Results on domain generalization task.
+TABLE III: Results on domain generalization task.
 Source Target
@@
 MMRL 72.03 64.47 49.17 51.20 77.53 60.59
+HiCroPL 71.22 64.33 49.47 50.79 77.15 60.44
+BIP 70.80 64.40 50.00 50.30 77.30 60.50
 ANPrompt 71.13 64.63 49.13 50.37 77.47 60.40
-Domain Generalization
+IV-E Domain Generalization
+We compare the performance of our ANPrompt with that of nine previous methods on cross-domain data sets.
+As shown in TABLE III, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability.
+Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
+IV-F Few-shot Learning
+We evaluate the discriminating ability of the model under limited supervision by changing the number of shots in each category (K = 1, 2, 4, 8, 16).
+The evaluation includes five baseline methods and ANPrompt proposed by us, such as CLIP, CoOp, CoCoOp, Prograd and Maple.
+As shown in Figure 6, ANPrompt is always superior to the comparison method in all lens settings, which shows the robustness of ANPrompt.
+Specifically, compared with the suboptimal method Maple, ANPrompt achieved an average performance improvement of 2.76%, 2.29%, 1.99%, 0.92% and 0.65% on 1, 2, 4, 8 and 16 shots on 11 data sets, respectively.
+In terms of a single data set, ANPrompt showed the best accuracy on 8 of 11 datasets.
+For example, in 16shots, ANPrompt achieved 74.27% value on DTD dataset, which was 2.94% higher than Maple.
+Compared with Maple, UCF101 is improved by 1.37%.
+In addition, ANPrompt’s performance on other datasets is still competitive.
+This shows that ANPrompt effectively enhances the robustness under weak semantic disturbance by injecting weak semantic disturbance into the alignment process and prompt representation.
+Figure 6: Few-shot classification accuracy across multiple datasets under varying shots. ANPrompt consistently outperforms existing methods.
 We compare the performance of our ANPrompt with that of six previous methods on cross-domain data sets.
-As shown in Table 3, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability.
+As shown in TABLE III, the average performance of ANPrompt on four data sets exceeds that of most comparison methods, showing its strong cross-domain generalization ability.
 Specifically, ANPrompt performs best on Imagenet-V2 dataset, and it also shows considerable performance on other target domain datasets.
-Ablation Experiments
-Componment Ablation
+IV-G Ablation Study
+IV-G1 Component ablation
 We conduct an ablation study to evaluate the contribution of each component in ANPrompt, including the Weak noise Frozen Text feature (TextNoise), Weak Alignment Loss (WALoss), and Anti-noise Prompt injection.
-As shown in Table 4, introducing WALoss or Anti-Prompt individually yields consistent improvements over the baseline, with gains of +0.06 and +0.19 respectively.
+As shown in Table IV, introducing WALoss or Anti-Prompt individually yields consistent improvements over the baseline, with gains of +0.06 and +0.19 respectively.
 Notably, the addition of TextNoise leads to further gains when combined with either WALoss (+0.34) or Anti-Prompt (+0.40).
@@
 These results verify the complementary nature of each module and their collective effectiveness in enhancing robustness under noise.
-Table 4: The Effectiveness of different components in ANPrompt. The TextNoise stands for the Weak noise Frozen Text feature.
+TABLE IV: The Effectiveness of different components in ANPrompt. The TextNoise stands for the Weak noise Frozen Text feature.
 TextNoise WALoss Anti-Prompt HM
@@
 81.70(+0.79)
-Effect of Prompt Token Length.
+IV-G2 Effect of Prompt Token Length
 We study the effect of varying the number of prompt tokens $T$ on performance.
-As shown in Table 5, increasing $T$ from 1 to 5 gradually improves performance across base and novel classes, leading to a peak harmonic mean (HM) of 81.70% at $T=5$.
+As shown in Table V, increasing $T$ from 1 to 5 gradually improves performance across base and novel classes, leading to a peak harmonic mean (HM) of 81.70% at $T=5$.
 This suggests that moderate prompt length enhances generalization.
@@
 Therefore, setting $T=5$ offers the best trade-off between expressiveness and stability.
-Table 5: The length of Prompt Token.
+TABLE V: The length of Prompt Token.
 $T$ 1 2 3 4 5 6
@@
 HMHM 80.53 80.30 80.45 80.66 81.70 81.66
-Effect of weight of Weak Noise Frozen Text feature.
+IV-G3 Effect of weight of Weak Noise Frozen Text feature
 We examine the sensitivity of ANPrompt to the weighting factor applied to the Weak Noise Frozen Text feature.
-As shown in Table 6, setting the weight to a small value (0.001) yields the best harmonic mean (81.70), indicating effective integration of weak noise signals.
+As shown in Table VI, setting the weight to a small value (0.001) yields the best harmonic mean (81.70), indicating effective integration of weak noise signals.
 As the weight increases, performance on novel classes consistently degrades, leading to a gradual drop in HM.
@@
 These results highlight the necessity of a carefully balanced contribution from TextNoise.
-Table 6: The Effectiveness of different weight of Weak Noise Frozen Text feature.
+TABLE VI: The Effectiveness of different weight of Weak Noise Frozen Text feature.
 Weight Base Novel HM
@@
 1.0 86.02 75.51 80.43
-Conclusion
+IV-G4 Logits Distance of WALoss
+We conducted an ablation experiment on the consistency constraint of WALoss ($\mathcal{L}_{\text{WA}}$).
+As shown in Table VII, we show the performance results when using different consistency constraints (including cosine, L1 and MSE).
+As you can see, our method ANPrompt has achieved the best results.
+Specifically, the result of ANPrompt in base is 86.15%, which is 5.57% higher than that of traditional MSE.
+The result on new is 77.70%, which is 5.82% higher than the traditional MSE.
+On the HM, the result of ANPrompt is 81.70%, which is 1.07%, 5.72% and 0.22% higher than L1, MSE and cosine, respectively.
+These results show that kl loss is most conducive to learning consistency under the condition of weak semantic perturbation to improve robustness.
+TABLE VII: Performance comparison of different calculative methods on WALoss ($\mathcal{L}_{\text{WA}}$).
+Method Base New HM
+L1 86.10 75.81 80.63
+MSE 80.58 71.88 75.98
+Cosine 85.27 78.01 81.48
+Ours 86.15 77.70 81.70
+IV-G5 Ablation study on $\theta$
+As shown in Table VIII, the HM score exhibits a consistent trend with increasing values of $\theta$.
+Specifically, performance gradually improves from $\theta=0.1$ to $\theta=0.7$, indicating that the progressive incorporation of alignment signals—despite being accompanied by weak semantic noise—introduces a beneficial level of learning difficulty.
+This increased difficulty encourages the model to more effectively acquire downstream semantic knowledge and enhance robustness.
+The performance peak at $\theta=0.7$ (81.70%) suggests an optimal trade-off, where the challenge introduced by noise is sufficient to promote generalization without overwhelming the model.
+However, as $\theta$ increases beyond $0.7$, a performance decline is observed, implying that excessive emphasis on alignment suppresses robustness-related cues and diminishes the advantages conferred by controlled noise.
+These findings highlight the role of moderate noise as a regularizing signal that facilitates generalization, while excessive alignment leads to suboptimal performance.
+TABLE VIII: Comparison of different number of $\theta$.
+Method Base New HM
+0.1 83.73 75.75 79.54
+0.2 84.36 75.48 79.67
+0.3 84.78 75.23 79.72
+0.4 85.17 76.32 80.50
+0.5 85.67 76.48 80.82
+0.6 85.98 76.93 81.20
+0.7 86.15 77.70 81.70
+0.8 86.10 75.85 80.65
+0.9 85.69 75.53 80.29
+1.0 84.75 75.38 79.79
+IV-G6 Ablation on different generation methods of $\gamma$
+We conduct an ablation study to evaluate different strategies for computing the weight $\gamma$ of the weak alignment loss. Specifically, we compare the following three formulations:
+TABLE IX: Ablation on different generation methods of $\gamma$.
+Method Base New HM
+Log 86.28 75.66 80.62
+mean 68.59 60.43 64.25
+Softmax 86.29 75.89 80.76
+Ours 86.15 77.70 81.70
+  • Log.
+This strategy applies logarithmic smoothing to the standard deviation to mitigate large gradients.
+However, the logarithmic transformation compresses variation, reducing sensitivity to changes in prediction confidence:
+    $\gamma=\frac{1}{\log(\mathrm{std}(\ell_{R})+1.1)\cdot|\ell_{R}|}$ (10)
+  • Mean. This method normalizes the variance by the mean logit value. However, it is highly unstable when the mean is small, often resulting in erratic scaling:
+    $\gamma=\frac{1}{\left(\frac{\mathrm{std}(\ell_{R})}{\mathrm{mean}(\ell_{R})+\epsilon}\right)\cdot|\ell_{R}|}$ (11)
+  • Softmax-Entropy. This approach uses output entropy as a proxy for uncertainty. While conceptually sound, it is sensitive to class imbalance and tends to saturate under confident predictions:
+    $\gamma=\frac{1}{\mathcal{H}(\mathrm{softmax}(\ell_{R}))\cdot|\ell_{R}|}$ (12)
+    where entropy is defined as:
+    $\mathcal{H}(p)=-\sum p\log(p)$ (13)
+As shown in Table IX, our proposed variance-adaptive weighting strategy achieves the highest HM score (81.70%) among all competing methods, demonstrating a clear advantage in overall performance.
+Compared to the Log and Mean baselines, the former applies logarithmic smoothing to the standard deviation to suppress gradient fluctuations but inadvertently compresses the model’s responsiveness to semantic variations.
+The latter normalizes the standard deviation by the mean, yet becomes numerically unstable when the mean approaches zero, often resulting in erratic weight scaling.
+The Softmax-Entropy method attempts to use output entropy as a proxy for uncertainty.
+Although conceptually reasonable, it tends to saturate under high-confidence predictions, failing to capture fine-grained semantic perturbations.
+In contrast, our approach dynamically adjusts the alignment strength based on the dispersion of the logits, maintaining stability while enhancing the model’s sensitivity to subtle semantic shifts.
+Moreover, it exhibits reduced sensitivity to class imbalance and low-magnitude outputs, thereby improving its generalization capability.
+Overall, the results suggest that variance-adaptive weighting strikes a better balance between stability and sensitivity, significantly enhancing the model’s robustness and cross-category generalization under weak semantic perturbations.
+IV-G7 Ablation on different noise injection methods
+TABLE X: Performance comparison of different noise injection.
+Method Base New HM
+Synreplace 85.90 77.43 81.44
+mask 85.99 77.41 81.47
+Shuffle 85.90 77.16 81.29
+Drop 85.97 76.94 81.21
+Ours 86.15 77.70 81.70
+Table X compares the performance of different noise injection methods, evaluated on Base accuracy, New class accuracy, and Harmonic Mean (HM).
+The methods tested include Synonym Replacement (Synreplace), Masking, Shuffling, Dropping, and our proposed method (Ours).
+Among the traditional methods, Synonym Replacement achieves the lowest performance, with a New class accuracy of 77.43% and an HM of 81.44%.
+Masking and Shuffling show slightly better results, with New class accuracies of 77.41% and 77.16%, and HMs of 81.47% and 81.29%, respectively.
+Dropping yields a New class accuracy of 76.94% and an HM of 81.21%.
+In contrast, our method (Ours) outperforms all others, with a Base accuracy of 86.15%, a New class accuracy of 77.70%, and the highest HM of 81.70%.
+This highlights the effectiveness of our approach in improving robustness to weak semantic noise while maintaining strong generalization to new categories.nOur method achieves this by integrating weak semantic perturbations into the prompt tokens and enforcing logits-level consistency via Weak Alignment Loss (WALoss).
+This design allows the model to handle noise-induced variations without sacrificing semantic fidelity, providing superior robustness compared to traditional noise injection techniques.
+IV-G8 Ablation on injection layers
+TABLE XI: Ablation on injection layers.
+Layers Base New HM
+$\{$1-3}\} 85.57 72.64 78.58
+$\{$1-6}\} 85.71 73.97 79.41
+$\{$1-9}\} 85.93 74.96 80.07
+$\{$1-12}\} 86.02 73.98 79.55
+$\{$3-6}\} 85.72 75.49 80.28
+$\{$3-9}\} 85.96 75.20 80.22
+$\{$3-12}\} 85.94 75.25 80.24
+$\{$6-9}\} 85.85 76.72 81.03
+$\{$6-12}\} 86.15 77.70 81.70
+$\{$9-12}\} 85.74 75.54 80.20
+As shown in the injection layers ablation results in Table XI, the layer position at which Anti-noise Prompts are injected has a significant impact on model performance.
+When prompts are injected into shallower layers (e.g., layers 1–3 or 1–6), the model achieves reasonable performance on Base classes but exhibits weaker generalization on New classes, along with a relatively lower Harmonic Mean (HM).
+As the injection layers move progressively deeper (e.g., layers 6–9 or 6–12), performance on both Base and New classes improves steadily, reaching its optimum when prompts are injected into layers 6–12, where the HM reaches 81.70%.
+This phenomenon may be attributed to the hierarchical nature of semantic information captured at different depths in vision-language models.
+Shallow layers typically extract local, low-level visual features, while deeper layers tend to model global, high-level semantic representations.
+Injecting Anti-noise Prompts into deeper layers (e.g., layers 6–12) allows for more effective introduction of mild perturbations at the semantic level, enabling the model to learn robustness to semantic variations in high-level feature spaces.
+As a result, the model maintains its discriminative ability on Base classes while significantly improving generalization to New classes.
+In contrast, injecting prompts into shallower layers (e.g., 1–3) may fail to adequately integrate semantic-level perturbations, whereas injection into even deeper layers (e.g., 9–12) could introduce excessive noise interference, leading to performance degradation.
+Therefore, selecting layers 6–12 as the injection interval achieves an optimal balance between semantic expressiveness and noise robustness.
+IV-H Grad-CAM Visualization
+Figure 7: Comparison of Grad-CAM visualizations.
+From left to right: original image, standard CLIP, CLIP with noisy text, and our ANPrompt with noisy text.
+The standard CLIP model is susceptible to textual noise, leading to inaccurate and dispersed attention.
+In contrast, our ANPrompt model effectively mitigates noise interference, maintaining concentrated and semantically precise heatmaps on the target objects.
+To evaluate the noise resistance of our model, we conducted a Grad-CAM gradcam visualization to compare the attention responses of the pre-trained CLIP model and ANPrompt under weak semantic noise.
+As shown in Figure  7, when the text input contains weak semantic perturbations, the attention maps of the CLIP model become significantly dispersed, with some focus on background or semantically irrelevant regions, indicating that its visual representations are susceptible to noise interference.
+In contrast, under the same noisy conditions, our ANPrompt model consistently maintains its attention on the key regions of the target object, demonstrating superior semantic robustness.
+This capability is notably evident on fine-grained datasets such as EuroSAT and FGVCAircraft, where ANPrompt accurately pinpoints discriminative local features, further validating its stable recognition performance in complex semantic scenarios.
+This observed robustness is attributed to the explicitly designed weak semantic noise training mechanism and the Noise-Resistant Visual Prompt Prototype (NRVPP) integrated into the ANPrompt framework.
+By clustering weak noise features and incorporating them into the prompt representations, our approach exposes the model to and helps it adapt to semantic variations during training, thereby enhancing its inherent resistance to noise.
+Concurrently, the proposed Weak Alignment Loss (WALoss) imposes a consistency constraint at the logits level, guiding the model to effectively filter out noise during inference and firmly anchor its predictions to the core semantic concepts.
+Consequently, ANPrompt suppresses semantic drift at the feature level while reinforcing discriminative power at the decision level, ultimately leading to more stable and interpretable visual attention.
+V CONCLUSION
 We proposed ANPrompt, a robust prompt tuning framework that explicitly models weak semantic perturbations to improve generalization in vision-language models.
@@
 Experimental results on 11 benchmarks demonstrate its clear advantages over existing methods, highlighting the effectiveness of incorporating weak semantic perturbation noise into prompt tuning.
-Acknowledgments
-This work is supported by the National Natural Science Foundation of China (62376106).